OpenStudy (anonymous):

Fine the perimeter of triangle A B C with vertices (-3,6), (-3,2), and (3,2). round your answer to two decimal places.

OpenStudy (anonymous):

Use the distance formula. In this case, since it is a right triangle, AB = 4, BC = 6. Using Pythagorean theorem, AC = sqrt [4^2 + 6^2] = sqrt (52) = 7.21 Finally, p = 4 + 6 + 7.21 = 17.21